OpenStudy (anonymous):

Multiply (x+5)(x+5)(x+5)= ?

OpenStudy (cwrw238):

first multiply (x + 5)(x + 5) can you do this?

OpenStudy (anonymous):

I put (x+5)(x+5) on my calculator and it gave 225

OpenStudy (anonymous):

So I don't need to put on my calculator (x+5)*(x+5)?

OpenStudy (anonymous):

Doing the three it gave 3375

OpenStudy (cwrw238):

no this is algebra and you use whats known as the Distributive Law first multiply the second parenthesis by x then by +5: (x + 5)(x + 5) = x(x + 5) + 5(x + 5) = x*x + x*5 + 5*x + 5*5 = x^2 + 10x + 25 are you familiar with algebra?

OpenStudy (anonymous):

Yea, Damn I need to do distribute property?

OpenStudy (cwrw238):

yes then second step is (x + 5)(x^2 + 10x + 25) = x(x^2 + 10x + 25) + 5(x^2 + 10x + 25)

OpenStudy (cwrw238):

and simplify as I did before

OpenStudy (anonymous):

Thanks bro.

OpenStudy (cwrw238):

yw